Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by checkley75(3666)

3C+G=1.75 OR G=1.75-3C
2C+4G=2.00 NOW SUBSTITUTE (1.75-3C) FOR G IN THIS EQUATION & SOLVE FOR C.
2C+4(1.75-3C)=2.00
2C+6.00-12C=2.00
-10C=2.00-6.00
-10C=-4.00
C=-4.00/-10
C=.40 COST OF THE GUM.
3C+.40=1.75
3C=1.75-40
3C-1.35
C=1.35/3
C=.45 COST OF THE CHOCOLATE.